5)Solve the following linear programming problem using two phase method. 

Minimize z = -3x1 + x2 - 2x3 

Subject to 

x1 +3x2 +x3 ≤5 

2x1 –x2 +x3 ≥2 

4x1 + 3x2 - 2x3 = 5 


x1, x2, x3 ≥ 0 




1
Expert's answer
2021-09-10T13:14:57-0400

SOLUTION


Changing the minimisation to maximization


Maximize z = 3x1 – x2 + 2x3

subject to

x1 + 3x2 + x3 ≤ 5

2x1 – x2 + x3 ≥ 2

4x1 + 3x2 - 2x3 = 5

x1, x2, x3 ≥ 0

Converting inequalities to equalities and introducing slack, surplus and artificial variables:

x1+ 3x2+ x3+ x4= 5


2x1– x2+ x3– x5+ A1= 2


4x1+ 3x2 - 2x3 + A2 = 5


x1, x2, x3, x4, x5, A1, A2 ≥ 0


Where:

A1 and A2 are artificial variables


x4 is a slack variable

x5 is a surplus variable



.

Phase 1 The objective function can be expressed as

Maximize 0x+ 0x2 + 0x3 + 0x4 + 0x5 + (–A1) + (–A2)

subject to

x1 + 3x2 + x3 + x4 = 5

2x1 – x2 + x3 – x5 + A1 = 2

4x1 + 3x2 - 2x3 + A2 = 5

x1, x2, x3, x4, x5, A1, A2 ≥ 0



Two Phase Method: Table 1




 c0 0 0 0 0 -1 -1 XB

cB Bv xxxxxAA2

0 x1 3 1 1 0 0 0 5

-1 A2 -1 1 0 -1 1 0 2

-1 A4 3 -2 0 0 0 1 5

zj–cj.  -6 -2 1 0 1 0 0 


Key column = x1 column

Key row = A1 row

Key element = 2

XB... solution values

A1 departs and xenters.


TABLE 2

CJ 0 0 0 0 0 -1 XB

CB BV X1 X2 X3 X4 X5 A2

0 X4 0 7/2 1/2 1 1/2 0 4

0 X1 1 -½ ½ 0 -½ 0 1

-1 A2 0 5 -4 0 2 1 1

Zj- Cj 0 -5 4 0 -2 0


A2 departs and xenters.

Both the artificial variables have been removed from the basis.


Phase 2:



TABLE 3



 c3 -1 2 0 0

CB BV xxxxx5 XB

0 x0 0 33/10 1 -9/13 33/10

3 x1 1 0 1/10 0 -3/10 11/10

-1 x2 0 1 -4/5 0 2/5 1/5

Zj-Cj 0 0 -3/10 0 -13/10


TABLE 4


 c3 -1 2 0 0 

cBv xxxxX5 XB

0 x0 9/4 3/2 1 0 15/4

3 x1 1 3/4 -1/2 0 0 5/4

0 x0 5/2 -2 0 1 1/2

zj-cj.  0 13/4 -7/2 0 0 






Two Phase Method: Optimal Table


  cj 3 -1 2 0 0  xb

Bv x1 x2 x3 x4 x5

2 x3 0 3/2 1 2/3 0 5/2

3 x11 3/2 0 1/3 0 5/2

0 x5 0 11/2 0 4/3 1 11/2

zj-cj  0 17/2 0 7/3 0 


An optimal is x= 5/2, x2 = 0, x3 = 5/2. The associated optimal value of the objective function is z = 3 X (5/2) – 0 + 2 X (5/2) = 25/2.